A kid pushes with a 100 N force on a 20kg lawnmower. The handle is elevated 37 degrees to the horizontal. What is the normal force exerted on the mower by the ground?
1
Expert's answer
2018-11-12T15:57:09-0500

The Newton’s second law gives

0=mg-N-F sin⁡θ

So, the normal force

N=mg+F sin⁡〖θ=20×9.8+100 sin⁡〖37°=256 N〗 〗